# An Asymptotically Optimal Coordinate Descent Algorithm for Learning Bayesian Networks from Gaussian Models

### v1

@@ 3. A Coordinate Descent Algorithm for DAG Learning
In this section, we develop a cyclic coordinate descent approach to find a heuristic solution to problem (3). The coordinate descent solver is fast and can be scaled to large-scale problems. As we demonstrate in Section 4, it provably converges and produces an asymptotically optimal solution to (3). Given the quality of its estimates, the proposed coordinate descent algorithm can also be used as a warm start for the mixed-integer programming framework in [25] to obtain optimal solutions.
3.1. Parameter Update without Acyclicity Constraints
@@
(4) $\min\limits_{\Gamma_{uv}\in\mathbb{R}}g(\Gamma_{uv})\coloneqq\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}\left(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}\right)+\lambda^{2}\|\Gamma-\mathrm{diag}(\Gamma)\|_{\ell_{0}},$
@@
Proof.
@@
$\mathrm{tr}\left(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}\right)$ =$\sum\limits_{i=1}^{m}\Gamma_{ui}\left(\Gamma_{ui}\hat{\Sigma}_{uu}+\sum_{j\not=u}\Gamma_{ji}\hat{\Sigma}_{ju}\right)+\sum\limits_{k\neq u}\sum\limits_{i=1}^{m}\Gamma_{ki}\left(\Gamma_{ui}\hat{\Sigma}_{uk}+\sum_{j\neq u}\Gamma_{ji}\hat{\Sigma}_{jk}\right).$
We first consider $\Gamma_{uv}$ for $u\not=v$. The derivative of $g(\Gamma_{uv})$ with respect to $\Gamma_{uv}$ is:
$\frac{\partial g(\Gamma_{uv})}{\partial\Gamma_{uv}}=\frac{\partial\mathrm{tr}(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})}{\partial\Gamma_{uv}}=2\hat{\Sigma}_{uu}\Gamma_{uv}+\sum\limits_{j\not=u}\Gamma_{jv}\hat{\Sigma}_{ju}+\sum_{k\not=u}\Gamma_{kv}\hat{\Sigma}_{uk}=2\hat{\Sigma}_{uu}\Gamma_{uv}+A_{uv}.$
Setting ${\partial g(\Gamma_{uv})}/{\partial\Gamma_{uv}}=0$, and defining $\hat{\gamma}_{uv}:=-A_{uv}/{2\hat{\Sigma}_{uu}}$, we obtain
@@
The original objective function $g$ with $\ell_{0}$-norm is nonconvex and discontinuous. To find the optimal solution, we compare $g(\hat{\gamma}_{uv})$ with $g(0)$. Given that $g(\hat{\gamma}_{uv})$ represents the optimal objective value for any nonzero $\Gamma_{uv}$, comparing it with $g(0)$ allows us to determine the optimal solution. Note that $g(\hat{\gamma}_{uv})-g(0)=\hat{\gamma}_{uv}^{2}\hat{\Sigma}_{uu}+\hat{\gamma}_{uv}A_{uv}+\lambda^{2}$. Thus, $g(\hat{\gamma}_{uv})\leq g(0)$ is equivalent to $\lambda^{2}\leq A_{uv}^{2}/{4\hat{\Sigma}_{uu}}$.
@@
Setting ${\partial g(\Gamma_{uu})}/{\partial\Gamma_{uu}}=0$, we obtain: $\hat{\Gamma}_{uu}={-A_{uu}\allowbreak+(A_{uu}^{2}\allowbreak+16\hat{\Sigma}_{uu}})^{1/2}\allowbreak/{4\hat{\Sigma}_{uu}}.$
3.2. Accounting for Acyclicity and Full Algorithm Description
@@
Algorithm 1 fully describes our procedure. The input to our algorithm is the sample covariance $\hat{\Sigma}$, regularization parameter $\lambda\in\mathbb{R}_{+}$, a super-structure graph $E_{\mathrm{super}}$ that is a superset of edges that contains the true edges, and a positive integer $C$. We allow the user to restrict the set of possible edges to be within a user-specified super-structure set of edges $E_{\text{super}}$. A natural choice of the superstructure is the moral graph, which can be efficiently and accurately estimated via existing algorithms such as the graphical lasso [7] or neighborhood selection [15]. This superstructure could also be the complete graph if a reliable superstructure estimate is unavailable.
We start by initializing $\Gamma$ as the identity matrix. Then, for each pair of indices $u$ and $v$ ranging from 1 to $m$, we update $\Gamma_{uv}$ based on specific rules. If $u=v$ (a diagonal entry), we update it directly according to Proposition 1. Among the off-diagonal entries, we only update those within the superstructure. Specifically, if $u\neq v$, and $(u,v)$ is in the superstructure, we check if setting $\Gamma_{uv}$ to a nonzero value violates the acyclicity constraint. (We use the breadth-first search algorithm [e.g., 6, 9] to check for acyclicity.) If it does not, we update $\Gamma_{uv}$ as per Proposition 1; otherwise, we set $\Gamma_{uv}$ to 0. We refer to a full sequence of coordinate updates as a full loop. The loop is repeated until convergence, when the objective values no longer improve after a complete loop. We keep track of the support of $\Gamma$s encountered during the algorithm. When the occurrence count of a particular support of $\Gamma$s reaches a predefined threshold, $C$, a spacer step [4, 11] is initiated, during which we update every nonzero coordinate iteratively. Note that in the spacer step, we use $\hat{\gamma}_{uv}$, which is the optimal update without considering the sparsity penalty, i.e., we use $\lambda^{2}=0$. The use of spacer steps stabilizes the behavior of updates and ensures convergence. After finishing the spacer step, we reset the counter of the support of the current solution.
1:  Input: Sample covariance $\hat{\Sigma}$, regularization parameter $\lambda\in\mathbb{R}_{+}$, super-structure $E_{\mathrm{super}}$, positive integer $C$.
2:  Initialize: $\Gamma^{0}\leftarrow I$; $t\leftarrow 1$.
3:  while objective function $f(\Gamma^{t})$ continue decreasing do
4:     for $u=1$ to $m$ do
5:        $\Gamma^{t}_{uu}=\hat{\Gamma}_{uu}$, where $\hat{\Gamma}_{uu}$ is calculated from Proposition 1 using the recently updated $\Gamma^{t}$.
6:        for $v=1$ to $m$ such that $(u,v)\in E_{\mathrm{super}}$ do
7:           If $\Gamma^{t}_{uv}\not=0$ violates acyclicity constraints, set $\Gamma^{t}_{uv}=0$.
8:           If $\Gamma^{t}_{uv}\not=0$ would not violate acyclicity constraints, set $\Gamma^{t}_{uv}=\hat{\Gamma}_{uv}$.
9:           $t\leftarrow t+1$
10:           $\mathrm{Count}[\mathrm{support}(\Gamma^{t})]\leftarrow\mathrm{Count}[\mathrm{support}(\Gamma^{t})]+1$.
11:           if $\mathrm{Count}[\mathrm{support}(\Gamma^{t})]=Cm^{2}$ then
12:                 $\Gamma^{t+1}\leftarrow\mathrm{SpacerStep}(\Gamma^{t})$           (Algorithm 2)
   $\mathrm{Count}[\mathrm{support}(\Gamma^{t})]=0.$
   $t\leftarrow t+1$.
13:           end if
14:        end for
15:     end for
16:  end while
17:  Output: $\hat{\Gamma}\leftarrow\Gamma^{t}$ and the Markov equivalence class $\mathrm{MEC}(\mathcal{G}(\hat{\Gamma}-\mathrm{diag}(\hat{\Gamma})))$
@@
Algorithm 2 SpacerStep
4. Statistical and Optimality Guarantees
We provide statistical and optimality guarantees for our coordinate descent procedure (Algorithm 1). Specifically, we follow a similar proof strategy as [11] to show that Algorithm 1 converges. Remarkably, we also prove the surprising result that the objective value attained by our coordinate descent algorithm provably converges to the optimal objective value of (3). Finally, we build on these results and provide finite-sample statistical consistency guarantees. Throughout, we assume the super-structure $E_{\mathrm{super}}$ that is supplied as input to Algorithm 1 satisfies $E^{\star}\subseteq E_{\mathrm{super}}$ where $E^{\star}$ denotes the true edge set; see [25] for a discussion on how the graphical lasso can yield super-structures that satisfy this property with high probability.
4.1. Convergence and Optimality Guarantees
@@
Assumption 1 (Positive definite sample covariance).
The sample covariance matrix $\hat{\Sigma}$ is positive definite.
@@
Assumption 1 is satisfied almost surely if $n\geq m$ and the samples of the random vector $X$ are generated from an absolutely continuous distribution. Under this mild assumption, our coordinate descent algorithm provably converges, as shown next.
Theorem 2 (Convergence of Algorithm 1).
@@
Let $\{\Gamma^{t}\}_{t=1}^{\infty}$ be the sequence of estimates generated by Algorithm 1. Suppose that Assumption 1 holds. Then,
  1. (1)
    the sequence $\{\mathrm{support}(\Gamma^{t})\}_{t=1}^{\infty}$ stabilizes after a finite number of iterations; that is, there exists a positive integer $M$ and a support set $\hat{E}\subseteq\{(i,j):i,j=1,2,\dots,m\}$ such that $\mathrm{support}(\Gamma^{t})=\hat{E}$ for all $t\geq M$.
  2. (2)
@@
    the sequence $\{\Gamma^{t}\}_{t=1}^{\infty}$ converges to a matrix $\Gamma$ with $\mathrm{support}(\Gamma)=\hat{E}$.
The proof of Theorem 2 relies on the following definitions and lemmas, and it closely follows the approach outlined in [11]. With a slight abuse of notation, we let $\ell(\Gamma):=\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}_{n})$ to be the negative log-likelihood function associated with parameter $\Gamma\in\mathbb{R}^{m\times m}$.
Definition 3 (Coordinate-wise (CW) minimum [11]).
A connectivity matrix $\Gamma^{\mathrm{CW}}\in\mathbb{R}^{m\times m}$ of a DAG is the CW minimum of problem (3) if for every $(u,v),u,v=1,\ldots,m$, $\Gamma^{\mathrm{CW}}_{uv}$ is a minimizer of $g(\Gamma_{uv})$ with other coordinates of $\Gamma^{\mathrm{CW}}$ held fixed.
Lemma 3.
Proof.
By Assumption 1, $\ell(\Gamma)$ is strongly convex and thus bounded below, and so is $f(\Gamma)$. If $\Gamma^{j}$ is the result of a non-spacer step, then the inequality $f(\Gamma^{j})\leq f(\Gamma^{j-1})$ holds trivially. Similarly, we know that if $\Gamma^{j}$ results from a spacer step, then, $\ell(\Gamma^{j})\leq\ell(\Gamma^{j-1})$. Since a spacer step updates only coordinates on the support, it cannot increase the support size of $\Gamma^{j-1}$, i.e., $\|\Gamma^{j}-\mathrm{diag}(\Gamma^{j})\|_{\ell_{0}}\leq\|\Gamma^{j-1}-\mathrm{diag}(\Gamma^{j-1})\|_{\ell_{0}}$, thus $f(\Gamma^{j})\leq f(\Gamma^{j-1})$. Since $f(\Gamma^{j})$ is non-increasing and bounded below, it must converge. ∎
Lemma 4.
Proof.
By Algorithm 1, $\Gamma^{t}\in G:=\{\Gamma\in\mathbb{R}^{m\times m}\mid f(\Gamma)\leq f(\Gamma^{0})\}$. It suffices to show that the set $G$ is bounded. From Proposition 11.11 in [3], if the function $f$ is coercive, then the set $G$ is bounded. Since $f(\Gamma)\geq\ell(\Gamma)$ for every $\Gamma$, it suffices to show that the function $\ell$ is coercive. By Assumption 1, we have that the function $\ell$ is strongly convex. The lemma then follows from the classical result in convex analysis that strongly convex functions are coercive. ∎
The following lemma characterizes the limit points of Algorithm 1.
Lemma 5.
  1. (1)
  2. (2)
  3. (3)
Proof.
Part 1.) Since spacer steps optimize only over the coordinates in $\hat{E}$, no element outside $\hat{E}$ can be added to the support. Thus, for every $l\in L$ we have $\mathrm{support}(\Gamma^{l})\subseteq\hat{E}$. We next show that strict containment is not possible via contradiction. Suppose $\operatorname{Supp}(\Gamma^{l})\subsetneq\hat{E}$ occurs infinitely often, and consider some $l\in L$ where this occurs. By the spacer step of Algorithm 1, the previous iterate $\Gamma^{l-1}$ has support $\hat{E}$, implying $\left\|\Gamma^{l-1}\right\|_{0}-\left\|\Gamma^{l}\right\|_{0}\geq$ 1. Moreover, from the definition of the spacer step, we have $\ell(\Gamma^{l})\leq\ell(\Gamma^{l-1})$. Therefore, we get $f(\Gamma^{l-1})-f(\Gamma^{l})=\ell(\Gamma^{l-1})-\ell(\Gamma^{l})+\lambda^{2}(\|\Gamma^{l-1}\|_{0}-\|\Gamma^{l}\|_{0})\geq\lambda^{2}$. Thus, when $\mathrm{support}(\Gamma^{l})\subsetneq\hat{E}$ occurs, $f$ decreases by at least $\lambda^{2}$. Therefore, $\Gamma^{l}\subsetneq\hat{E}$ infinitely many times implies that $f(\Gamma)$ is not lower-bounded, which is a contradiction.
Part 2.) The proof follows the conventional procedure for establishing the convergence of cyclic coordinate descent (CD) [4, 11]. We obtain $\Gamma^{l}$ by updating every coordinate in $\hat{E}$ of $\Gamma^{l-1}$. Denote the intermediate steps as $\Gamma^{l,1},\ldots,\Gamma^{l,|\hat{E}|}$, where $\Gamma^{l,|\hat{E}|}=\Gamma^{l}$. We aim to show that the sequence $\{\Gamma^{l,|\hat{E}|}\}_{l\in L}$ converges to a point $\Gamma^{\mathrm{CW}}$, and similarly, other sequences $\{\Gamma^{l,i}\}_{l\in L},i=1,\ldots,|\hat{E}|-1$, also converge to $\Gamma^{\mathrm{CW}}$. By Lemma 4, since $\{\Gamma^{l,|\hat{E}|}\}_{l\in L}$ is a bounded sequence, there exists a converging subsequence $\{\Gamma^{l^{\prime},|\hat{E}|}\}_{l^{\prime}\in L^{\prime}}$ with a limit point $\Gamma^{\mathrm{CW}}$. Without loss of generality, we choose the subsequence satisfying $l^{\prime}>M$, $\forall l^{\prime}\in L^{\prime}$. From Part 1 of the lemma, $\{\Gamma^{l^{\prime},1}\}_{l^{\prime}\in L^{\prime}},\ldots,\{\Gamma^{l^{\prime},|\hat{E}|-1}\}_{l^{\prime}\in L^{\prime}}$ all have the same support $\hat{E}$. For $\{\Gamma^{l^{\prime},|\hat{E}|-1}\}_{l^{\prime}\in L^{\prime}}$, we have $f(\Gamma^{l^{\prime},|\hat{E}|-1})-f(\Gamma^{l^{\prime},|\hat{E}|})=\ell(\Gamma^{l^{\prime}|\hat{E}|-1})-\ell(\Gamma^{l^{\prime},|\hat{E}|})$. If the change from $\Gamma^{l^{\prime},|\hat{E}|-1}$ to $\Gamma^{l^{\prime},|\hat{E}|}$ is on a diagonal entry, say $\Gamma_{uu}$, then, after some algebra, we obtain
$\ell\left(\Gamma^{l^{\prime},|\hat{E}|-1}\right)-\ell\left(\Gamma^{l^{\prime},|\hat{E}|}\right)=\\ 2\left(-\log{\Gamma_{uu}^{l^{\prime},|\hat{E}|-1}}/{\Gamma_{uu}^{l^{\prime},|\hat{E}|}}+{\Gamma_{uu}^{l^{\prime},|\hat{E}|-1}}/{\Gamma_{uu}^{l^{\prime},|\hat{E}|}}-1\right)+\left(\Gamma_{uu}^{l^{\prime},|\hat{E}|-1}-\Gamma_{uu}^{l^{\prime},|\hat{E}|}\right)^{2}\hat{\Sigma}_{uu}.$
Since $a-1\geq\log(a)$ for $a\geq 0$, each of the two terms above is non-negative. From Lemma 3, as $l^{\prime}\to\infty$, $f(\Gamma^{l^{\prime},|\hat{E}|-1})-f(\Gamma^{l^{\prime},|\hat{E}|})$ or equivalently $\ell(\Gamma^{l^{\prime},|\hat{E}|-1})-\ell(\Gamma^{l^{\prime},|\hat{E}|})$ converges to 00 as $l^{\prime}\rightarrow\infty.$ Combining this with the fact that $\ell(\Gamma^{l^{\prime},|\hat{E}|-1})-\ell(\Gamma^{l^{\prime},|\hat{E}|})\geq 0$ and that each term in the equality for $\ell(\Gamma^{l^{\prime},|\hat{E}|-1})-\ell(\Gamma^{l^{\prime},|\hat{E}|})$ is non-negative, we conclude that $\Gamma^{l^{\prime},|\hat{E}|-1}$ must converge to $\Gamma^{l^{\prime},|\hat{E}|}$ as $l^{\prime}\to\infty$. Since $\Gamma^{l^{\prime},|\hat{E}|}$ converges to $\Gamma^{\mathrm{CW}}$, $\Gamma^{l^{\prime},|\hat{E}|-1}$ must also converge to $\Gamma^{\mathrm{CW}}$. Repeating a similar argument, we conclude that $\Gamma^{l^{\prime},j}$ converges to $\Gamma^{\mathrm{CW}}$ for all $j=1,2,\dots,|\hat{E}|$.
Part 3.) Consider any subsequence $\{\Gamma^{k}\}_{k\in K}$ such that $\mathrm{support}(\Gamma^{k})=\hat{E}$. We will show by contradiction that $\{\Gamma^{k}\}_{k\in K}$ must converge to $\Gamma^{\mathrm{CW}}$. Suppose $\{\Gamma^{k}\}_{k\in K}$ has a limit point $\hat{\Gamma}\not=\Gamma^{\mathrm{CW}}$. Then there exist a subsequence $\{\Gamma^{k^{\prime}}\}_{k^{\prime}\in K^{\prime}}$, with $K^{\prime}\subseteq K$, that converges to $\hat{\Gamma}$. Therefore, $\lim_{k^{\prime}\rightarrow\infty}f(\Gamma^{k^{\prime}})=\ell(\hat{\Gamma})+\lambda^{2}|\hat{E}|.$ From part 1 and part 2, we have that $\lim_{l^{\prime}\rightarrow\infty}f(\Gamma^{l^{\prime}})=\ell(\Gamma^{\mathrm{CW}})+\lambda^{2}|\hat{E}|.$ By Lemma 3, we have $\lim_{k^{\prime}\rightarrow\infty}f(\Gamma^{k^{\prime}})=\lim_{l^{\prime}\rightarrow\infty}f(\Gamma^{l^{\prime}})$. Thus, we conclude that $\ell(\hat{\Gamma})=\ell(\Gamma^{\mathrm{CW}})$, which contradicts the fact that $\Gamma^{\mathrm{CW}}$ is the unique minimizer of $\min_{\mathrm{support}(\Gamma)\subseteq\hat{E}}\ell(\Gamma)$. Therefore, we conclude that any subsequence with support $\hat{E}$ converges to $\Gamma^{\mathrm{CW}}$ as $k\rightarrow\infty$. ∎
Lemma 6.
Let $\Gamma$ be a limit point of $\{\Gamma^{k}\}_{k=1}^{\infty}$ with $\mathrm{support}(\Gamma)=\hat{E}$. Then we have $\mathrm{support}(\Gamma^{k})=\hat{E}$ for infinitely many $k$’s.
Proof.
We prove this result by contradiction. Assume that there are only finitely many $k$’s such that $\mathrm{support}(\Gamma^{k})=\hat{E}$. Since there are finitely many possible support sets, there is a support $E^{\prime}\not=\hat{E}$ and a subsequence $\{\Gamma^{k^{\prime}}\}$ of $\{\Gamma^{k}\}$ such that $\mathrm{support}(\Gamma^{k^{\prime}})=E^{\prime}$ for all $k^{\prime}$, and $\lim_{k^{\prime}\rightarrow\infty}\Gamma^{k^{\prime}}=\Gamma$. However, by Lemma 5, the subsequence converges to a minimizer $\Gamma^{\mathrm{CW}}$ with $\mathrm{support}(\Gamma^{\mathrm{CW}})=E^{\prime}$ and thus $\Gamma^{\mathrm{CW}}\not=\Gamma$. This is a contradiction. ∎
We are now ready to complete the proof of Theorem 2.
Proof of Theorem 2.
Let $\Gamma$ be a limit point of $\{\Gamma^{k}\}$ with the largest support size and denote its support by $\hat{E}$. By Lemma 6, there is a subsequence $\{\Gamma^{r}\}_{r\in R}$ of $\{\Gamma^{k}\}$ such that $\mathrm{support}(\Gamma^{r})=\hat{E},\forall r\in R$, and $\lim_{r\rightarrow\infty}\Gamma^{r}=\Gamma$. By Lemma 5, there exists an integer $M$ such that for every $r\geq M$ and $r+1$ is a spacer step, we have $\mathrm{support}(\Gamma^{r})=\mathrm{support}(\Gamma^{r+1})$. Without loss of generality, we choose the subsequence that $r>M,\forall r\in R$. We will demonstrate by contradiction that any coordinate $(u,v)$ in $\hat{E}$ cannot be dropped infinitely often in $\{\Gamma^{k}\}$. To this end, assume that $(u,v)\not\in\{\mathrm{support}(\Gamma^{r})\}_{r>M}$ infinitely often. Let $\{\Gamma^{r^{\prime}}\}_{r^{\prime}\in R^{\prime}}$, where $R^{\prime}\subseteq R$, be the subsequence with $\mathrm{support}(\Gamma^{r^{\prime}+1})=\hat{E}\setminus\{(u,v)\},\forall r^{\prime}\in R^{\prime}$. Since $r^{\prime}>M$ and the support has been changed, $r^{\prime}+1$ is not a spacer step. Therefore, using Proposition 1, we have $f(\Gamma^{r^{\prime}})-f(\Gamma^{r^{\prime}+1})\geq\lambda^{2}-{A^{2}_{uv}}/{4\hat{\Sigma}_{uu}}>0$. By Lemma 3, we have $\lim_{r^{\prime}\rightarrow\infty}f(\Gamma^{r^{\prime}})-f(\Gamma^{r^{\prime}+1})=0$. Thus, $\lambda^{2}={A^{2}_{uv}}/{4\hat{\Sigma}_{uu}}$, where $A_{uv}=\sum_{j\not=u}\Gamma^{r^{\prime}}_{jv}\hat{\Sigma}_{ju}+\sum_{k\not=u}\Gamma_{kv}^{r^{\prime}}\hat{\Sigma}_{uk}$. By Proposition 1, in step $r^{\prime}+1$, we have $|\Gamma^{r^{\prime}+1}_{uv}|={\lambda}/{\sqrt{\hat{\Sigma}_{uu}}}>0$, which contradicts the definition of $\{\Gamma^{r^{\prime}}\}_{r^{\prime}\in R^{\prime}}$. Therefore, no coordinate in $\hat{E}$ can be dropped infinitely often. Moreover, no coordinate can be added to $\hat{E}$ infinitely often as $\hat{E}$ is the largest support. As a result, the support converges to $\hat{E}$. With stabilized support $\hat{E}$, by Lemma 5, we have that $\{\Gamma^{k}\}$ converges to the limit $\Gamma^{\mathrm{CW}}$ with support $\hat{E}$. From Algorithm 1 and Proposition 1, we have $\Gamma_{uv}$ is a minimizer of $f(\Gamma_{uv})$ with respect to the coordinate $(u,v)$ and others fixed. Therefore, $\Gamma^{\mathrm{CW}}$ is the CW minimum. ∎

### v3

@@ 3 A Coordinate Descent Algorithm for DAG Learning
In this section, we develop a cyclic coordinate descent approach to find a heuristic solution to problem (3). The coordinate descent solver is fast and can be scaled to large-scale problems. As we demonstrate in Section 4, it provably converges and produces an asymptotically optimal solution to (3). Given the quality of its estimates, the proposed coordinate descent algorithm can also be used as a warm start for the mixed-integer programming framework in Xu et al. (2024) to obtain optimal solutions.
3.1 Parameter update without acyclicity constraints
@@
$\min\limits_{\Gamma_{uv}\in\mathbb{R}}g(\Gamma_{uv})\coloneqq\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}\left(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}\right)+\lambda^{2}\|\Gamma-\mathrm{diag}(\Gamma)\|_{\ell_{0}},$ (4)
@@
Proof The original objective function $g$ with $\ell_{0}$-norm is nonconvex and discontinuous. To find the optimal solution, we compare $g(0)$ with $g(\hat{\gamma}_{uv})$ for some $\hat{\gamma}_{uv}\not=0$.
@@
[…]
$\frac{\partial\mathrm{tr}(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})}{\partial\Gamma_{uv}}=2\hat{\Sigma}_{uu}\Gamma_{uv}+\sum\limits_{j\not=u}\Gamma_{jv}\hat{\Sigma}_{ju}+\sum_{k\not=u}\Gamma_{kv}\hat{\Sigma}_{uk}=2\hat{\Sigma}_{uu}\Gamma_{uv}+A_{uv}.$
Setting the derivative to zero, and defining $\hat{\gamma}_{uv}:=-A_{uv}/{(2\hat{\Sigma}_{uu})}$, we obtain
@@
Furthermore, we observe that $\frac{\partial^{2}}{\partial\Gamma_{uv}^{2}}\mathrm{tr}(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})=2\hat{\Sigma}_{uu}>0$. This positive second derivative implies that optimizing $g(\Gamma_{uv})$ for any nonzero off-diagonal entry reduces to solving a convex optimization problem.
Given that $g(\hat{\gamma}_{uv})$ represents the optimal objective value for any nonzero $\Gamma_{uv}$, comparing it with $g(0)$ allows us to determine the optimal solution. Note that $g(\hat{\gamma}_{uv})-g(0)=\hat{\gamma}_{uv}^{2}\hat{\Sigma}_{uu}+\hat{\gamma}_{uv}A_{uv}+\lambda^{2}$. Thus, $g(\hat{\gamma}_{uv})\leq g(0)$ is equivalent to $\lambda^{2}\leq A_{uv}^{2}/{(4\hat{\Sigma}_{uu})}$.
@@
Setting ${\partial g(\Gamma_{uu})}/{\partial\Gamma_{uu}}=0$, we obtain: $\hat{\Gamma}_{uu}=({-A_{uu}\allowbreak+(A_{uu}^{2}\allowbreak+16\hat{\Sigma}_{uu}})^{1/2})\allowbreak/{(4\hat{\Sigma}_{uu})}.$
3.2 Accounting for acyclicity and full algorithm description
@@
Algorithm 1 fully describes our procedure. The input to our algorithm is the sample covariance $\hat{\Sigma}$, regularization parameter $\lambda\in\mathbb{R}_{+}$, a super-structure graph $E_{\mathrm{super}}$ that is a superset of edges that contains the true edges, a positive integer $C$, and a topological ordering $O$ that is a permutation of $\{1,2,\dots,m\}$. We allow the user to restrict the set of possible edges to be within a user-specified super-structure set of edges $E_{\text{super}}$. A natural choice of the superstructure is the moral graph, which can be efficiently and accurately estimated via existing algorithms such as the graphical lasso (Friedman et al., 2007), neighborhood selection (Meinshausen and Buhlmann, 2006), or $\ell_{0}$-regularized pseudo-likelihood-based estimator (Behdin et al., 2023). This superstructure could also be the complete graph if a reliable superstructure estimate is unavailable. The input topological ordering $O$ governs the order in which the coordinates of $\Gamma$ are updated in our coordinate descent algorithm (see Remark 4 for more discussion).
[…]
Remark 4
The effect of the input ordering $O$ is that Algorithm 1 iteratively updates the coordinates of $\Gamma$ in a particular order: updating row by row in the order given by O{O}, and within each row, according to O{O}. In Section 4, we show that regardless of the input ordering $O$, as well as the initialization $\Gamma_{\mathrm{init}}$, Algorithm 1 converges, and is asymptotically optimal (matches the optimal objective value as the sample size tends to infinity). However, while different orderings result in similar final objective values, the associated Markov equivalence classes can vary significantly. Thus, a given ordering or initialization may result in a Markov equivalence class that is far from the population one. In Section 5, we discuss this challenge, and advocate for using an existing algorithm to find a suitable ordering, and use this ordering as input to Algorithm 1. Finally, we note that the input topological ordering $O$ is simply an initialization to Algorithm 1. Indeed, the topological ordering corresponding to the output $\hat{\Gamma}$ can in general be different than $O$. Finally, we note that the default initialization $\Gamma^{\mathrm{init}}=I$ tends to perform well in practice and is comparable to other initialization values, especially when a good ordering $O$ is selected; see Section 6.
@@
1:Input: Sample covariance $\hat{\Sigma}$, regularization parameter $\lambda\in\mathbb{R}_{+}$, super-structure $E_{\mathrm{super}}$, positive integer $C$, any topological ordering $O$ that is a permutation of $\{1,2,\dots,m\}$, initial matrix $\Gamma^{\mathrm{init}}$ (default: identity matrix).
2:Initialize: $\Gamma^{0}\leftarrow\Gamma^{\mathrm{init}}$ ; $t\leftarrow 1$;
[…]
4:while objective function $f(\Gamma^{t})$ continue decreasing do
5:  for $u=1$ to $m$ do
6:   $\Gamma^{t}_{uu}=\hat{\Gamma}_{uu}$, where $\hat{\Gamma}_{uu}$ is calculated from Proposition 3 using the recently updated $\Gamma^{t}$.
7:   for $v=1$ to $m$ such that $(u,v)\in E_{\mathrm{super}}$ do
8:    If $\Gamma^{t}_{uv}\not=0$ violates acyclicity constraints, set $\Gamma^{t}_{uv}=0$.
9:    If $\Gamma^{t}_{uv}\not=0$ would not violate acyclicity constraints, set $\Gamma^{t}_{uv}=\hat{\Gamma}_{uv}$.
10:    $t\leftarrow t+1$.
11:    $\mathrm{Count}[\mathrm{support}(\Gamma^{t})]\leftarrow\mathrm{Count}[\mathrm{support}(\Gamma^{t})]+1$.
12:    if $\mathrm{Count}[\mathrm{support}(\Gamma^{t})]=Cm^{2}$ then
13:      $\Gamma^{t+1}\leftarrow\mathrm{SpacerStep}(\Gamma^{t})$.      (Algorithm 2) $\mathrm{Count}[\mathrm{support}(\Gamma^{t})]=0.$
14:     $t\leftarrow t+1$.
15:    end if
16:   end for
17:  end for
18:end while
[…]
Algorithm 2 SpacerStep
@@
4 Convergence and Optimality Guarantees
We provide convergence and optimality guarantees for our coordinate descent procedure (Algorithm 1). Specifically, we follow a similar proof strategy as Hazimeh and Mazumder (2020) to show that Algorithm 1 converges. Remarkably, we also prove the surprising result that the objective value attained by our coordinate descent algorithm provably converges to the optimal objective value of (3). Throughout, we assume the super-structure $E_{\mathrm{super}}$ that is supplied as input to Algorithm 1 satisfies $E^{\star}\subseteq E_{\mathrm{super}}$ where $E^{\star}$ denotes the true edge set; see Xu et al. (2024) for a discussion on how the graphical lasso can yield super-structures that satisfy this property with high probability.
Our analysis relies on the notion of coordinate-wise minimum (Hazimeh and Mazumder, 2020), which represent solutions Algorithm 1 obtains under different input orderings $O$.
Definition 5
(Coordinate-wise (CW) minimum) A connectivity matrix $\Gamma^{\mathrm{CW}}\in\mathbb{R}^{m\times m}$ of a DAG is the CW minimum of problem (3) if for every $(u,v),u,v=1,\ldots,m$, $\Gamma^{\mathrm{CW}}_{uv}$ is a minimizer of $g(\Gamma_{uv})$ with other coordinates of $\Gamma^{\mathrm{CW}}$ held fixed.
4.1 Convergence
@@
Assumption 1
(Positive definite sample covariance) The sample covariance matrix $\hat{\Sigma}$ is positive definite.
@@
Assumption 1 is satisfied almost surely if $n\geq m$ and the samples of the random vector $X$ are generated from an absolutely continuous distribution. Under this mild assumption, our coordinate descent algorithm provably converges to a coordinate-wise minimum.
Theorem 6
@@
(Convergence of Algorithm 1) Let $\{\Gamma^{t}\}_{t=1}^{\infty}$ be the sequence of estimates generated by Algorithm 1. Suppose that Assumption 1 holds. Then, for input topological ordering $O$ or initial point $\Gamma^{\mathrm{init}}$ to Algorithm 1:
  1. 1.
[…]
  2. 2.
@@
    the sequence $\{\Gamma^{t}\}_{t=1}^{\infty}$ converges to a coordinate-wise minimum $\Gamma$ with $\mathrm{support}(\Gamma)=\hat{E}$.
The proof of Theorem 6 is provided in Appendix A and adapts the analysis in Hazimeh and Mazumder (2020), which was originally developed for coordinate descent in variable selection, to our problem setting.
4.2 Optimality guarantees